OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5760 — Backbone in Lille
Scheduled Maintenance Report for Web Cloud
Completed
We will implement the new optical loop
on the first loop of Lille with NRA security.

The NRA that will be impacted by this maintenance:

LIB59 SEC59 FAC59 BOI59 BLE59 PIL59

The beginning of the maintenance is scheduled at midnight in the
night between 07/09/11 to 08/09/11

The maximum cut is 6 o'clock.

Changes in maintenance will
complete this task.




Update(s):

Date: 2011-09-08 23:45:04 UTC
A new task has been opened:
http://status.ovh.co.uk/?do=details&id=1786&PHPSESSID=7a01e6a51d515c519ac4340f13417347

Date: 2011-09-08 06:57:01 UTC
Blueberry NRA is UP now. We are reporting the switch of the optical buckle. Some points to check in order to proceed to the complete implementation of the buckle in one night.


Date: 2011-09-08 06:54:53 UTC
Actually no sites are cut. Optical power measures took a long time. However, we are going to cut the blueberry site for some time in order to validate a first optical bypass. We will keep you informed of the evolution in a working task.
Posted Sep 07, 2011 - 14:38 UTC